Demand over the next four months are 450, 500, 350, and 400 respectively. Regular time production costs $10 per item, and we can make a maximum of 400 items in a given month in regular time. Overtime production costs $16 per item, and as many as 200 items may be produced in overtime each month (in addition to the regular time production). Keeping an item in inventory incurs a cost of $1 per period; backorders cost $3 per period per item. If we follow a level strategy of production, what would be the production quantity each month?
Answer #1

Total Demand over 4 months = 450+500+350+400 = 1700

Average monthly Demand = 1700/4 = 425 units

Regular time production capacity is 400 units per month

So 400 units are produced in regular time every month

50 units each are produced in first two months using overtime.

50 units are backordered in second month, which will be supplied in third month.

Using level strategy, production Quantity in each month = 400 units in regular time